# Improving Graph Neural Network Training, Defense and Hypergraph Partitioning via Adversarial Robustness Evaluation

### v9

@@ Abstract
Graph Neural Networks (GNNs) are a highly effective neural network architecture for processing graph-structured data. Unlike traditional neural networks that rely solely on the features of the data as input, GNNs leverage both the graph structure, which represents the relationships between data points, and the feature matrix of the data to optimize their feature representation. This unique capability enables GNNs to achieve superior performance across various tasks. However, it also makes GNNs more susceptible to noise and adversarial attacks from both the graph structure and data features, which can significantly increase the training difficulty and degrade their performance. Similarly, a hypergraph is a highly complex structure, and partitioning a hypergraph is a challenging task. This paper leverages spectral adversarial robustness evaluation to effectively address key challenges in complex-graph algorithms. By using spectral adversarial robustness evaluation to distinguish robust nodes from non-robust ones and treating them differently, we propose a training-set construction strategy that improves the training quality of GNNs. In addition, we develop algorithms to enhance both the adversarial robustness of GNNs and the performance of hypergraph partitioning. Experimental results show that this series of methods is highly effective.
@@ 1 Introduction
We also provide a efficient high performance hypergraph partitioning method based on spectral adversarial robustness evaluation and demonstrate its effectiveness on benchmark VLSI data sets.
@@ 3.2.4 Results
4 Improving Hypergraph partitioning
In this section, we propose a robustness-aware hypergraph partitioning framework. We first identify and extract a robust sub-hypergraph consisting of vertices that are empirically less sensitive to perturbations, and perform the partitioning algorithm only on this reduced structure. The resulting partition labels are then propagated to the remaining non-robust vertices by exploiting the underlying graph topology. In this way, our method reduces the effective size of the hypergraph (in terms of the number of vertices) while largely avoiding the degradation in solution quality that typically arises from changes to the hypergraph topology. The main contributions of this paper are summarized as follows:
  • We propose a novel robustness-aware hypergraph partitioning method that leverages adversarial robustness evaluation to distinguish robust nodes from non-robust nodes and treats them differently during the partitioning process.
  • We propose a highly novel label propagation scheme for propagating partition labels within hypergraphs.
  • We conduct extensive experiments on diverse hypergraph instances extracted from realistic VLSI design problems and demonstrate that our approach achieves clearly improved performance over existing hypergraph coarsening methods.
4.1 Method
Our method is built on the key idea of treating nodes that are robust to perturbations separately from those that are vulnerable. For nodes that are less susceptible to perturbations, we directly apply a hypergraph partitioning algorithm to obtain their partitions. In contrast, for nodes that are more easily affected by such perturbations, we do not rely on hypergraph spectral partitioning; instead, we infer their partition membership based on their relationships with the robust nodes and propagate the partitioning results from the robust subset to these vulnerable nodes. Specifically, as shown in Figure 2, our method consists of the following five phases.
@@ 4.1 Method
4.1.1 Phase 1: Hypergraph Spectral Embedding
We first perform a spectral embedding of the hypergraph by mapping it into the feature space spanned by the $K$ eigenvectors of its Laplacian matrix corresponding to the smallest eigenvalue magnitudes.
4.1.2 Phase 2: Manifold Capturing
Data manifolds are commonly modeled using graphs, with k-nearest neighbor (k-NN) graphs being one of the most prevalent choices. In this stage, we construct a k-NN graph to capture the underlying manifold structure of the spectrally embedded hypergraph. A standard k-NN construction procedure has a time complexity of $O(|N|^{2})$, where $|N|$ denotes the number of nodes, which becomes prohibitive for large-scale datasets. To address this issue, we adopt an approximate k-NN graph construction algorithm [25], which extends a probabilistic skip-list–like structure and reduces the complexity to $O(|N|\log|N|)$, making it far more suitable for large-scale data processing.
4.1.3 Phase 3: Adversarial Robustness Evaluation
Let $L_{X}$ and $L_{Y}$ denote the Laplacian matrices associated with the hypergraph $H$ and the $k$-NN graph constructed from the spectrally embedded hypergraph in phase 2, respectively. Following the theoretical framework in [24], we exploit the dominant generalized eigenvalues of $L_{Y}^{+}L_{X}$ and their corresponding eigenvectors to characterize the sensitivity of each sample to noise under a given model. In particular, we compute a weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$ to obtain a spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ is the number of nodes. The matrix $V_{s}$ is defined as
@@ 4.1.3 Phase 3: Adversarial Robustness Evaluation
4.2 Phase 4: Sub-Hypergraph Partitioning
For the nodes in the hypergraph $H$ we first sort them in descending order according to their Spade scores and select the bottom portion to form the robust node set. Using these robust nodes together with the hyperedges that connect them in the original hypergraph, we construct a robust sub-hypergraph. We then apply a standard hypergraph partitioning algorithm to this sub-hypergraph to obtain a partition of the robust nodes.
4.3 Label Propagation from Robust to Non-Robust Nodes
The nodes that are not included in the robust node set are referred to as non-robust nodes. Since these nodes are vulnerable to adversarial perturbations and attacks that affect hypergraph spectral partitioning, we do not directly apply the hypergraph spectral partitioning algorithm to determine their partitions. Instead, we infer their partition membership based on their relationships with the robust nodes, effectively propagating the partitioning result from the robust subset to the non-robust nodes. Specifically, for each non-robust node, we identify the robust node with which it co-occurs in hyperedges in $H$ most frequently, and assign the non-robust node to the same partition as that robust node.
@@ 4.3 Label Propagation from Robust to Non-Robust Nodes
Algorithm 1 Robust Hypergraph Partitioning
Input: A hypergraph $H=(V,E)$
Output: Partitioned hypergraph.
1: Extract the adjacency matrix AA from the input hypergraph $H$.
2: Compute the Laplacian matrix $\mathbf{L_{H}}$ for the graph $H$.
3: Compute the eigenvectors $u_{1}$$u_{k}$ that correspond to the bottom k nonzero eigenvalues of AA;
4: Construct $U\in\mathbb{R}^{n\times k}$, with $k$ eigenvectors of AA stored as columns;
5: Construct a $k$-nearest neighbor graph $G=(V,E)$ from $U$.
6: Compute the Laplacian matrix $\mathbf{L_{U}}$ for the graph $U$.
7: Compute the SPADE score of each node.
8: Form a robust node set from nodes with small SPADE scores, while constructing a non-robust node set from nodes with large SPADE scores.
9: Construct a sub-hypergraph $S$ consisting of the robust node set and all hyperedges in $H$ that connect only these robust nodes.
10: For each non-robust node, count the number of hyperedges in $H$ where it appears together with each robust node.
11: Apply standard hypergraph spectral partitioning on $S$.
12: For each non-robust node, identify the robust node with which it co-occurs in hyperedges in $H$ most frequently, and assign the non-robust node to the partition containing that robust node.
13: Return the partitioned hypergraph.
@@ 4.4 Experiments
We perform extensive experiments to evaluate the performance of the proposed spectral hypergraph coarsening method on a suite of VLSI-related benchmark datasets, from “ibm01” to “ibm04”Datasets are available at: https://vlsicad.ucsd.edu/UCLAWeb/cheese/ispd98.html.. To ensure a fair comparison, we also apply several existing hypergraph coarsening techniques to the same set of benchmarks and compare their results against ours. For brevity, we refer to the star and clique models as S and C, respectively. The experimental results are shown in Table 3. It can be seen that our method consistently outperforms all competing methods across all four datasets.
Table 3: The average local conductance ($\Phi$) comparison between our methods and other methods
Hypergraph V E Spectral S Spectral C Metis S Metis C hMETIS HyperSF Ours
ibm01 12752 14111 0.94 0.81 0.84 0.78 0.65 0.44 0.28
ibm02 19601 19584 0.94 0.84 0.83 0.79 0.69 0.52 0.43
ibm03 23136 27401 0.95 0.86 0.85 0.79 0.67 0.48 0.41
ibm04 27507 31970 0.95 0.84 0.84 0.78 0.68 0.47 0.40
@@ 5 Strengthen GNN Defenses via Adversarial Robustness Evaluation
The proposed method for improving the robustness of GNNs, which consists of five key phases. Phase (1) learns a latent node representation: we train a GNN and use the activations of its hidden layer as embeddings for the original feature matrix. In Phase (2), we construct $k$-NN graphs to capture the underlying manifold structure of the embedded feature space. In Phase (3), we quantify the robustness of each node by jointly leveraging information from the input graph and the learned output manifold. In Phase (4), we select a subset of robust nodes, extract the induced subgraph consisting only of them from the original graph, and apply a GNN to classify these robust nodes. In Phase (5), we perform label propagation based on the connectivity patterns between robust and non-robust nodes. In the rest of this section, we provide a detailed description of these five phases.
@@ 5.1 Method
5.1.1 Phase 1: Learning GNN Latent Representations
[24] introduced a framework that analyzes the adversarial robustness of individual edges by constructing a bijective distance mapping between the manifolds of a model’s input space and output space.
In our setting, we first train a GNN on the original feature matrix together with the given input graph to obtain a well-fitted model. After training, this GNN is used to map the input features into a latent representation space. Since the last layer of a GNN is typically tailored to the downstream task (for example, a softmax layer in the case of multi-class node classification), we use the activations of the penultimate hidden layer as the embedding of each node, rather than the final task-specific outputs.
5.1.2 Phase 2: Capturing the Data Manifold
A common way to approximate the underlying data manifold is to represent it with a graph, where nodes correspond to data points and edges reflect local neighborhood relationships. Among such constructions, $k$-nearest neighbor ($k$-NN) graphs are particularly popular. In this phase, we construct a $k$-NN graph on top of the embedded feature matrix obtained from the GNN, so as to model the manifold structure in the latent space.
The standard $k$-NN graph construction procedure has a time complexity of $O(|N|^{2})$, where $|N|$ denotes the number of nodes, which becomes prohibitive for large-scale datasets. To address this scalability issue, we adopt an approximate $k$-NN graph construction method [25]. This approach builds on an extension of probabilistic skip-list–like structures to significantly reduce the computational cost, achieving a time complexity of $O(|N|\log|N|)$ and thus making it suitable for large data regimes.
5.1.3 Phase 3: Spectral Adversarial Robustness Evaluation
Let $L_{X}$ and $L_{Y}$ denote the Laplacian matrices of the original input graph and the $k$-NN graph constructed in the embedded feature space, respectively. Following the theoretical framework in [24], we leverage the leading generalized eigenvalues of $L_{Y}^{+}L_{X}$ together with their associated eigenvectors to estimate how sensitive each sample is to noise under a specified model. Specifically, we form a weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$, which serves as a spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ denotes the number of nodes. The matrix $V_{s}$ is defined as
@@ 5.1.3 Phase 3: Spectral Adversarial Robustness Evaluation
5.1.4 Phase 4: Sub-graph Classification
For the nodes in the original graph, we first rank them in descending order by their Spade scores and then take the lowest-scoring fraction to define a robust node set. Using these robust nodes and the edges among them inherited from the original graph, we build a robust subgraph. Subsequently, we train a GNN on this robust subgraph to perform node classification for the robust nodes.
5.1.5 Phase 5: Label Propagation from Robust to Non-Robust Nodes
@@ 5.2 Experiments
We evaluate robustness using an independent (non-cumulative) white-box greedy edge-deletion attack. For each target test node $v$, we start from a fresh copy of the clean graph $G$ and perform at most $T$ deletion steps. At step $t$, we enumerate all incident edges $(v,u)$, temporarily remove each candidate edge, and run a forward pass to measure the logit of the true class $y_{v}$ at node $v$. We then permanently delete the edge that yields the largest decrease in this true-class logit. The attack stops early if $v$ becomes misclassified, if no incident edges remain, or if removing any remaining edge no longer decreases the true-class logit. This procedure is applied independently to each target node (i.e., perturbations are not shared across targets). We select as effective targets those test nodes that are correctly classified on the clean graph by both the baseline and our method, and report the post-attack accuracy on these targets (equivalently, $1-\mathrm{ASR}$). The experimental results are shown in Table 4. It can be seen that our method is very effective.
Table 4: Post-attack accuracy on effective target nodes under the independent white-box greedy edge-deletion attack (Cora).
Model Effective Targets Baseline Acc. Ours Acc.
GAT 200 0.7450 0.7900
GCN 200 0.7300 0.7650

### v11

# Improving Graph Neural Network Training, Defense and Hypergraph Clustering via Adversarial Robustness Evaluation

@@ Abstract
Graph Neural Networks (GNNs) are a highly effective neural network architecture for processing graph-structured data. Unlike traditional neural networks that rely solely on the features of the data as input, GNNs leverage both the graph structure, which represents the relationships between data points, and the feature matrix of the data to optimize their feature representation. This unique capability enables GNNs to achieve superior performance across various tasks. However, it also makes GNNs more susceptible to noise and adversarial attacks from both the graph structure and data features, which can significantly increase the training difficulty and degrade their performance. Similarly, a hypergraph is a highly complex structure, and partitioning a hypergraph is a challenging task. This paper leverages spectral adversarial robustness evaluation to effectively address key challenges in complex-graph algorithms. By using spectral adversarial robustness evaluation to distinguish robust nodes from non-robust ones and treating them differently, we propose a training-set construction strategy that improves the training quality of GNNs. In addition, we develop algorithms to enhance both the adversarial robustness of GNNs and the performance of hypergraph clustering. Experimental results show that this series of methods is highly effective.
@@ 1 Introduction
We also provide a efficient high performance hypergraph clustering method based on spectral adversarial robustness evaluation and demonstrate its effectiveness on benchmark hypergraphs.
@@ 3.2.4 Results
4 Improving Hypergraph Clustering
Our method is built on the key idea of treating nodes that are robust to noise and adversarial perturbations separately from those that are vulnerable. For nodes that are less susceptible to noise and perturbations, we directly apply a standard spectral hypergraph clustering algorithm to obtain their labels. In contrast, for nodes that are more easily affected by such noise and perturbations, we do not rely on the clustering algorithm; instead, we infer their labels based on their relationships with the robust nodes and propagate the clustering results from the robust subset to these vulnerable nodes. Specifically, as shown in Figure 2, our method consists of the following five phases.
@@ 4 Improving Hypergraph Clustering
4.1 Phase 1: Graph Spectral Embedding
To obtain a spectral embedding, the hypergraph is first converted into a clique expansion graph via clique expansion. The $k$ eigenvectors of the Laplacian of the clique expansion graph associated with the $k$ smallest-magnitude eigenvalues are then used to embed its vertices into a $k$-dimensional spectral space.
4.2 Phase 2: Manifold Capturing
Data manifolds are commonly modeled using graphs, with k-nearest neighbor (k-NN) graphs being one of the most prevalent choices. In this phase, we construct a k-NN graph to capture the underlying manifold structure of the spectrally embedded space.
4.3 Phase 3: Adversarial Robustness Evaluation
Let $L_{X}$ and $L_{Y}$ denote the Laplacian matrices associated with the clique expansion graph $G_{X}$ and the $k$-NN graph constructed from the spectrally embedded space in phase 2, respectively. Following the theoretical framework in [23], we exploit the dominant generalized eigenvalues of $L_{Y}^{+}L_{X}$ and their corresponding eigenvectors to characterize the sensitivity of each sample to adversarial perturbations under a given model. In particular, we compute a weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$ to obtain a spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ is the number of nodes. The matrix $V_{s}$ is defined as
@@ 4.3 Phase 3: Adversarial Robustness Evaluation
4.4 Phase 4: Subgraph Extraction
For the nodes in the clique expansion graph $G_{X}$ we first sort them in descending order according to their Spade scores and select the bottom portion to form the robust node set. Using these robust nodes together with the edges that connect them in the $G_{X}$, we construct a robust subgraph.
4.5 Subgraph Clustering and Label Propagation
A standard spectral graph clustering algorithm is applied to the robust subgraph to obtain cluster labels for robust nodes. Nodes outside the subgraph are referred to as non-robust nodes. Since non-robust nodes are more vulnerable to adversarial perturbations that can degrade spectral clustering, their labels are not determined by directly clustering the full graph. Instead, their labels are inferred by propagating the robust clustering results through $G_{X}$. Specifically, each non-robust node is assigned to the cluster that receives the largest total edge weight from its robust neighbors in $G_{X}$.
@@ 4.5 Subgraph Clustering and Label Propagation
Algorithm 1 Robust Spectral Hypergraph Clustering
Input: A hypergraph $\mathcal{G}=(V,\mathcal{E})$, number of clusters $k$, robust ratio $\rho$.
Output: Vertex clusters $\{C_{1},\dots,C_{k}\}$.
1: Construct the incidence matrix $H\in\mathbb{R}^{n\times m}$ from $\mathcal{G}$.
2: Form the clique expansion adjacency matrix $A_{X}\leftarrow HH^{\top}$; set $A_{ii}\leftarrow 0$.
3: Compute the normalized Laplacian $L_{X}\leftarrow I-D_{X}^{-1/2}A_{X}D_{X}^{-1/2}$.
4: Compute a spectral embedding $U_{e}$ of $G_{X}$ using $K_{\text{embed}}$ eigenvectors of $L_{X}$ with smallest magnitudes; row-normalize $U_{e}$.
5: Construct a k-NN graph $G_{Y}$ from the rows of $U_{e}$ and compute its normalized Laplacian $L_{Y}$.
6: Compute the coupled Laplacian matrix $M\leftarrow L_{Y}^{+}L_{X}$ using a numerically stabilized pseudoinverse.
7: Compute the top-$s$ dominant eigenpairs $(\zeta_{i},v_{i})$ of $M$, and construct the weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{n\times s}$ where column $i$ is $v_{i}\sqrt{\zeta_{i}}$.
8: For each node $p$, compute its SPADE score: $\text{Spade}(p)=\frac{1}{|N_{X}(p)|}\sum_{q\in N_{X}(p)}\|V_{s}(p)-V_{s}(q)\|_{2}^{2}$, where $N_{X}(p)$ is the neighbor set of $p$ in $G_{X}$.
9: Sort nodes by SPADE score in descending order and select the lowest-$\rho n$ nodes as the robust node set $R$.
10: Extract the induced robust subgraph $G_{R}$ from $G_{X}$ using nodes in $R$.
11: Perform standard spectral graph clustering on $G_{R}$ to obtain clusters for robust nodes $\{C_{1}^{R},\dots,C_{k}^{R}\}$.
12: Initialize labels for all nodes: $\ell(v)\leftarrow 0$; assign robust nodes labels according to the clustering result in $G_{R}$.
13: For each non-robust node $u\notin R$, infer its cluster by weighted majority vote from its robust neighbors in $G_{X}$ using edge weights in $A_{X}$.
14: Return final clusters $\{C_{1},\dots,C_{k}\}$.
@@ 4.6 Experiments
In this section, we conduct extensive experiments to evaluate the performance of the proposed robust spectral hypergraph clustering algorithm.
4.7 Data
We evaluate the proposed method on the following three benchmark hypergraphs. The key statistics are summarized in Table 3.
  • Zoo: A small categorical benchmark where vertices represent animals and hyperedges are formed by grouping animals sharing the same attribute value. The dataset contains seven ground-truth classes.
  • Tic-Tac-Toe Endgame: A dataset of terminal tic-tac-toe board states with nine categorical position attributes. Hyperedges connect states that share the same symbol at a given position, and the labels indicate two outcome classes.
  • Car Evaluation: A categorical dataset for car quality assessment. We construct a hypergraph by creating one hyperedge for each attribute–value group, and use the four provided labels for evaluation.
Table 3: Statistics of hypergraphs used in our experiments.
Hypergraphs Vertices Hyperedges Total Incidences Clusters
Zoo 101 35 1615 7
Tic-Tac-Toe Endgame 958 27 8622 2
Car Evaluation 1728 21 10368 4
4.8 Evaluation Metric
To evaluate clustering performance, we adopt the commonly used measure: clustering accuracy (ACC). The metric is computed between the cluster assignments produced by the clustering methods and the ground-truth labels provided in the data sets. The clustering accuracy is defined as:
$ACC=\frac{\sum\limits_{j=1}^{n}{\delta{(y_{i},map(c_{i}))}}}{{n}},$ (2)
where $n$ denotes the total number of data instances. $y_{i}$ represents the ground-truth label of the $i$-th instance, and $C_{i}$ is the label assigned by the clustering algorithm. $\delta(x,y)$ is the indicator function: $\delta(x,y)$=1 if $x=y$, and $\delta(x,y)$=0 otherwise. The function $map(\bullet)$ permutes cluster indices to match ground-truth labels, which can be implemented via the Hungarian algorithm [papadimitriou1998combinatorial]. Larger values of $ACC$ correspond to higher clustering quality.
4.9 Experimental Results
As shown in Table 4, our method consistently outperforms standard hypergraph clustering in terms of clustering accuracy (ACC) across all three benchmarks. Specifically, ACC increases from 62.38% to 80.20% on Zoo, from 51.67% to 59.39% on Tic-Tac-Toe Endgame, and from 34.43% to 46.88% on Car Evaluation, yielding improvements of 17.82, 7.72, and 12.45 percentage points, respectively. These gains demonstrate the effectiveness of our approach.
Table 4: Clustering accuracy (ACC) Results (%).
Hypergraphs Standard Hypergraph Clustering Ours
Zoo 62.38 80.20
Tic-Tac-Toe Endgame 51.67 59.39
Car Evaluation 34.43 46.88
@@ 5 Strengthen GNN Defenses via Adversarial Robustness Evaluation
In this section, we propose a method to defend GNNs using adversarially robust nodes. Specifically, we first conduct a spectral adversarial robustness evaluation and partition the nodes in the graph into two groups: adversarially robust nodes and adversarially non-robust nodes. We then extract a subgraph consisting of the adversarially robust nodes from the original graph and apply a GNN to classify these nodes. Finally, we propagate the labels from robust nodes to non-robust nodes based on the graph topology. Experiments on benchmark datasets with popular GNN architectures, evaluated under both Metattack and Nettack attacks, demonstrate that the proposed method yields highly effective defense against adversarial attacks.
@@ 5.1 Method
As shown in Figure 3, our method consists of six steps, which are detailed in the rest of this section.
Figure 3: Overview of the proposed method.
5.1.1 Step 1: Learning Latent Representations
Machine learning models can be viewed as transformation engines that convert inputs into outputs by successively reshaping representations. For example, deep neural networks compress and refine raw feature vectors through multiple nonlinear layers. Likewise, methods such as Support Vector Machines (SVM) and Support Vector Clustering (SVC) employ kernel functions to implicitly lift data from the original feature space into a higher-dimensional one [25]. Adversarial attacks seek to exploit this mechanism by inducing failures in the input–output transformation, causing the model to produce incorrect predictions. Therefore, we first run the model in its standard (benign) setting to obtain the model-induced transformation of the input feature space, i.e., the resulting latent representations.
5.1.2 Step 2: Capturing Manifolds
The data manifold is commonly modeled using graphs. Among the many graph construction strategies, the $k$-nearest neighbor ($k$-NN) graph is one of the most widely used and is well suited for manifold approximation. Accordingly, in this phase we construct $k$-NN graphs to represent the manifolds of both the original input feature matrix and the embedded feature matrix obtained in Phase 1.
5.1.3 Step 3: Spectral Adversarial Robustness Evaluation
Let $L_{X}$ and $L_{Y}$ denote the Laplacian matrices associated with the k-NN graphs built in Phase 2 from the input feature matrix and the embedded feature matrix, respectively. Following the theoretical framework in [23], we leverage the leading generalized eigenvalues of $L_{Y}^{+}L_{X}$ together with their associated eigenvectors to estimate how sensitive each sample is to noise under a specified model. Specifically, we form a weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$, which serves as a spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ denotes the number of nodes. The matrix $V_{s}$ is defined as
@@ 5.1.3 Step 3: Spectral Adversarial Robustness Evaluation
5.1.4 Step 4: Subgraph Extraction
For the nodes in the original graph, we first rank them in descending order by their Spade scores and then take the lowest-scoring fraction to define a robust node set. Using these robust nodes and the edges among them inherited from the original graph, we build a robust subgraph.
5.1.5 Step 5: Subgraph Classification
Subsequently, we train a GNN on this robust subgraph to perform node classification for the robust nodes.
5.1.6 Step 6: Label Propagation from Robust to Non-Robust Nodes
@@ 5.2 Experiments
5.2.1 Data Sets
We evaluate the proposed method on two standard citation network benchmarks, where nodes represent documents, edges encode citation links, and each node is associated with a sparse bag-of-words feature vector and a class label. A brief summary of the dataset characteristics is provided below.
  • Cora: A citation network of 2,708 nodes and 5,429 edges, where each node is represented by a 1,433-dimensional bag-of-words feature vector and assigned to one of 7 research categories.
  • Citeseer: A citation network of 3,327 nodes and 4,732 edges with 3,703-dimensional bag-of-words features. Compared to Cora, Citeseer is sparser and has a more imbalanced label distribution over its 6 classes, thereby posing a greater challenge for graph-based classification models.
Following ProGNN [26] and the implementation in the DeepRobust library[27], we adopt the predefined splits provided by Dataset(..., setting="prognn"). For each dataset, we operate on the largest connected component of the citation graph and use 10% of the nodes as labeled training data, 10% as validation data, and the remaining 80% as test data. This setting corresponds to a semi-supervised scenario with scarce labels and a relatively large test set, and is widely used in prior work on adversarial attacks and defenses on graphs [26, 27].
5.2.2 GNN Models
We use the following three popular GNN models for our experiments:
  • Generalized PageRank GNN (GPR-GNN): An architecture that combines adaptive generalized PageRank (GPR) scheme with GNNs, effectively mitigating feature over-smoothing[28].
5.3 Detailed Model Implementation and Parameter Settings
  • GCN: We adopt a two-layer GCN model, where the first GCN layer maps the input node features to an 8-dimensional hidden representation, followed by a ReLU activation, and the second GCN layer projects the hidden representations to the output class space.
  • GAT: We employ a two-layer GAT model. The first graph attention layer applies multi-head attention (with 4 heads in our experiments) to transform the input node features into an 8-dimensional hidden representation per head, whose outputs are concatenated, passed through an ELU activation. The second attention layer uses a single head without concatenation to map the hidden representations to the output class space.
  • GPR-GNN: We adopt a GPR-GNN architecture that consists of a two-layer MLP followed by a generalized PageRank (GPR) propagation module. The MLP first maps the input node features to a 16-dimensional hidden representation using a linear layer with ReLU activation, and then uses a second linear layer to project the hidden representation to the output class logits. On top of these logits, we apply a GPR-based propagation module with $K=10$ propagation steps over the GCN-normalized adjacency matrix with self-loops. This module aggregates information from multiple hop distances using a set of learnable propagation coefficients, which are initialized according to a Personalized PageRank scheme with the restart probability set to $0.1$.
For all experiments, we select 5% of the nodes as non-robust nodes. The number of neighbors in the k-NN graph is set to $k=10$, and we use 10 eigenvectors for the spectral embedding in the adversarial robustness evaluation.
5.3.1 Experimental Results
The experimental results of Metattack and Nettack are shown in Table 5 and Table 6, respectively.
Table 5: Classification accuracy (%) of three GNN models on two datasets, with and without our defense, under clean graph and Metattack. The last column reports the defense gain on attacked graphs.
Model Dataset Clean (Orig) Metattack (Orig) Clean (Ours) Metattack (Ours) Defense Gain
GCN Cora 82.49 56.34 77.11 66.45 10.11
GCN Citeseer 70.79 61.02 70.62 66.17 5.15
GAT Cora 82.29 42.25 79.33 66.50 24.25
GAT Citeseer 62.09 49.59 60.96 57.35 7.76
GPR-GNN Cora 80.03 69.37 81.79 76.61 7.24
GPR-GNN Citeseer 71.03 56.93 69.31 69.02 12.09
Table 6: Classification accuracy (%) of three GNN models on two datasets, with and without our defense, under clean graph and Nettack. The last column reports the defense gain on attacked graphs.
Model Dataset Clean (Orig) Nettack (Orig) Clean (Ours) Nettack (Ours) Defense Gain
GCN Cora 82.49 62.42 77.11 74.75 12.33
GCN Citeseer 64.04 53.61 65.17 63.57 9.96
GAT Cora 82.49 69.97 82.04 81.89 11.92
GAT Citeseer 56.64 41.59 57.41 52.13 10.54
GPR-GNN Cora 74.65 60.31 68.56 67.81 7.50
GPR-GNN Citeseer 67.95 58.53 68.36 66.94 8.41